# A Training-Free Length Extrapolation Approach for LLMs: Greedy Attention Logit Interpolation (GALI)

### v1

@@ 3.2 Attention Logit Interpolation
When calculating attention scores, we approximate attention logits for interpolated positions using local linear interpolation based on interpolated position IDs, while introducing Gaussian noise that scales with relative position intervals. Unlike NTK or YaRN, our methods avoid outlier attention logits from the position embeddings (Chen et al., 2023b). RoPE’s trigonometric functions, while effective within pretrained relative position intervals, can produce extreme values during interpolation, leading to abnormal attention logits. To mitigate this, we bypass position embeddings and directly approximate attention logits. Building on the monotonic trends and oscillatory behavior observed in RoPE, as shown in Figure 1, we hypothesize that for two tokens with an interpolated relative distance (e.g., $7.r$), their attention logits—excluding oscillatory effects—should lie between $g(7)$ and $g(8)$, with their exact value determined by $r$. To achieve this, we compute attention logits for each QK pair with interpolated relative position intervals using local linear interpolation and simulate oscillatory behavior by adding Gaussian noise, whose variance increases with relative position intervals. This method eliminates the outliers identified in the PI study and produces interpolated attention logits for new relative position intervals in a training-free and effective manner.
Specifically, given two token embeddings $\bm{x}_{m},\bm{x}_{n}\in\mathbb{R}^{l}$ as query and key corresponding to position $m$ and $n$. Note here the $m$ and $n$ can be interpolated float position ids. Let $r=m-n$, when ${r}{\%}1=0$ we compute their attention logits directly which means their relative position interval was pretrained. For the case ${r}{\%}1>0$, which means that their relative positional interval is interpolated, we then compute the attention logit $a(\bm{x}_{m},\bm{x}_{n},r)$ using the formula:
$a(\bm{x}_{m},\bm{x}_{n},r)=a(\bm{x}_{m},\bm{x}_{n},\lfloor{r}\rfloor)-[a(\bm{x}_{m},\bm{x}_{n},\lfloor{r}\rfloor))-$ (3)
$a(\bm{x}_{m},\bm{x}_{n},\lceil{r}\rceil]*({r}{\%}1)+\mathcal{N}(0,{\frac{r}{L_{tr}}}^{2})$
[…]
@@ 4 Experiments
We evaluate GALI on Llama3-8B-ins models across two task categories: real-world long-context tasks and long-context language modeling tasks. For comparison, we implement all published training-free length extrapolation methods, including NTK(LocalLLaMA, 2023b), Dyn-NTK(LocalLLaMA, 2023a), YARN(Peng et al., 2023), SelfExtend(Jin et al., 2024), and ChunkLlama(An et al., 2024b). The following sections describe the experimental setup for each task. Appendix B details data statistics.
4.1 Emperiments Setup
Real-world long-context task: We evaluate GALI on two widely used long-context benchmarks, LongBench(Bai et al., 2024) and L-Eval(An et al., 2024a). LongBench is a bilingual long-context understanding benchmark with 21 datasets crossing 6 tasks, including Single-Document QA, Multi-Document QA, Summarization, Few-shot Learning, Synthetic Task, and Code Completion. We use 16 English datasets from LongBench. L-Eval is a comprehensive long-context evaluation suite with 20 sub-tasks, 508 long documents, and over 2000 human-labeled query-response pairs. It consists of closed-ended and open-ended task groups. We focus on closed-ended groups, which assess reasoning and understanding over a long-context. For consistency, we follow the official task prompt templates and truncation strategies from the respective benchmarks.
Long-context language modeling task: To evaluate GALI’s long-context language modeling capabilities, we use PG19(Rae et al., 2019), an open-vocabulary language modeling benchmark derived from Project Gutenberg. PG19 consists of over 28,000 books published before 1919, covering diverse genres and writing styles. We use its test split, which includes 100 samples.
@@ 4.1 Emperiments Setup
Table 1: Performance comparison across different backbone LLMs and training-free length extrapolation methods. The best result in each experiment is bolded. Results marked with * are reported by LongBench (Jin et al., 2024). The number following each method denotes the target context window size (e.g., 16k represents 16 × 1024 tokens). ”Original” refers to evaluations conducted using the backbone model in the left column. Additional results using the Llama2-7B-Chat-4K backbone are provided in Appendix D.1.
Backbone models and baseline methods: We use Llama3-8b-ins-4k (Llama3-4k) and Llama3-8b-ins-8k (Llama3-8k) as backbone models, where the number following each model indicates its initial context window size. We obtain Llama3-4k backbone via modifying its max_position_embedding parameter. We use shorter-versions of Llama3-8b-ins over other LLMs with shorter training context windows like LLama2 since it cannot fully understand all pretrained positional intervals, which limits GALI’s effectiveness in practice. The effective understanding range of LLMs is shorter than their maximum training window, as evidenced in (Jin et al., 2024; Hsieh et al., 2024). For the baseline methods, we compare with NTK(LocalLLaMA, 2023b), Dyn-NTK(LocalLLaMA, 2023a), YaRN(Peng et al., 2023) using huggingface implementation and SelfExtend(Jin et al., 2024), ChunkLlama(An et al., 2024b) with their official implementation. They are all of the training-free length extrapolation methods up to now. The implementation details of these methods can be found in Appendix C.
@@ 4.2 Real-World long-context Task Results
The LongBench results (Table 1) highlight GALI’s strong average performance on the Llama3-8b-ins backbone series, surpassing both the 4k and 8k backbone models and other methods. GALI excels in question answering (QA) and few-shot learning tasks but showed weaker performance in summarization, synthetic, and code-related tasks. While extending the context window to 32k improves performance on Llama3-8k compared to 16k, it remains less effective than using 16k on Llama3-4k. This consistent trend in Figure 3(a) highlights limitations in extrapolation performance, which we analyze below.
Firstly, LLMs interpret positional intervals differently within their training context window, as noted in (Hsieh et al., 2024). This explains why using a 16k context window on Llama3-4k outperformed using the same window on Llama3-8k. Since LLMs are trained via next-token prediction, smaller positional intervals are more trained and thus better understood. The difference between models lies in the range of positional intervals they can fully comprehend based on their training context window. As a result, all methods except ChunkLlama perform better with a 16k context window on the Llama3-4k than with the same context window on the Llama3-8k. Secondly, despite all methods aiming for training-free length extrapolation, they employ fundamentally different approaches. GALI, NTK, Dyn-NTK, and YaRN focus on generalizing positional intervals, creating smaller but more usable intervals from pretrained ones. Among these, GALI generates only the minimal number of new intervals required, maximizing the reliance on pretrained intervals. Other methods predefine a target context window and fix the number and size of new intervals. Dyn-NTK only decides whether to generate new intervals based on input length without minimizing their quantity. SE and ChunkLlama do not create new positional intervals but instead redistribute inputs within existing intervals through input rearrangement. This leads to differences in practice: GALI maps inputs to the full range of the positional intervals of the training context window, while NTK, Dyn-NTK, YaRN and SelfExtend map inputs to predefined positional interval ranges based on the target context window (e.g. in SelfExtend, a larger context window leads to a larger group size). ChunkLlama determines interval ranges via hyperparameters such as chunk size and local window. The performance gap when using 16k and 32k context windows on Llama3-8k reflects these differences. Figure 6(b) shows that most LongBench dataset lengths are below 16k when using the Llama3 tokenizer. However, NTK, Dyn-NTK, YaRN and SelfExtend still benefit significantly from expanding to 32k, particularly on HotpotQA and Musique [^2] SelfExtend is highly sensitive to its hyperparameters, as mentioned in its GitHub implementation. Hence, our reproduced results using the Llama3-8k with 32k context window are significantly higher than those reported in their paper.. This is because most of the data in HotpotQA and Musique is concentrated below 16k. A 32k context window can map them to the positional interval range [0–4096), while a 16k one maps them to the range [0–8192) (For SelfExtend, this range is not fixed and depends on the specific local window and group size. However, it typically falls within a smaller positional interval range.). In contrast, GALI’s results for data shorter than 16k remain nearly unchanged, while ChunkLlama exhibits slight fluctuations. Overall, these methods that do not rely on predefined target context windows show limited improvement since most dataset lengths remain below 16K. However, on NarrativeQA—the longest dataset—GALI shows a substantial improvement from 25.88 to 28.63, outperforming all other methods.
@@
Table 2: Performance comparison across different backbone LLMs and training-free length extrapolation methods. The best result in each experiment is bolded. The GSM, QuALITY, and TOEFL were excluded from this comparison because their sequence lengths remain below 8192 tokens when using the Llama3 tokenizer, making them unsuitable for long-context evaluation. Results for these three datasets, along with those for all datasets using the Llama2-7B-Chat-4K (Llama2-4k) backbone, are provided in Appendix D.1.
@@
The L-Eval results further support our analysis. As shown in Table 2, GALI achieved the highest average performance, except when using a 32k context window on the Llama3-8k. Consistent with LongBench, where 32k on Llama3-8k performed slightly better than 16k but remained inferior to 16k on Llama3-4k, aligning with our previous observations. Since Llama3 better understands closer positional intervals, other methods benefit from mapping text to a smaller range of positional intervals, giving GALI a relative disadvantage at 32k on Llama3-8k. However, GALI still achieved the second-best performance. To investigate further, we tested GALI with a 32k context window on Llama3-4k, forcing it to operate exclusively within the [0, 4096) range of positional intervals. Under these conditions, GALI once again achieved the best results. Figure 3(b) shows the overall performance trends. Additionally, it is worth noting that when using the Llama3 tokenizer, GSM, QuALITY, and TOEFL lengths are all below 8192, where SE, YaRN, and NTK outperform the backbone models as shown in Appendix D.1, further validating our findings.
Experiments on LongBench and L-Eval show that training-free length extrapolation methods benefit from mapping text to smaller relative positional intervals, compensating for models’ uneven positional understanding. GALI, however, minimizes new intervals while utilizing the full range of pretrained intervals, which can be disadvantageous in some cases. As models improve, GALI’s direct attention logit estimation enhances context comprehension, delivering superior stability without requiring hyperparameter tuning.
@@
Table 3: Performance of Diverse Models on the PG19 Dataset Across Varying Context Window Sizes.
@@ 4.3 Long Language Modeling Task Results
The language modeling results are shown in Table 3. Due to OOM, we cannot get ChunkLlama’s PPL results when setting the maximum position embedding to 32768. Except for NTK and DYN-NTK, all methods maintained a stable PPL without exploding. While low PPL does not guarantee better real-world task performance, an exploding PPL is a clear indicator of performance degradation in downstream tasks. Notably, GALI achieved the second-lowest PPL, demonstrating superior stability in length extrapolation. We tested PPL using a 16k contest window with Llama2-4k backbone. Please refer to the Appendix D.2.
@@
Figure 4: Differences in attention score metrics and row-wise entropy across methods compared to the original LLM. (a) represents the attention score matrix differences. The values show the sum of the absolute differences between the attention scores of the length interpolation methods and the model trained on its training context window. (b) and (c) show row-wise entropy differences. 2K indicates methods using [0, 2048) positional intervals, while 4k shows the methods using positional intervals within [0, 4096). The figures are generated using a sample from NarrativeQA with a prefill length of 8091 tokens. More details on the attention score distributions can be found in Appendix D.3.
@@ 4.4 Attention Distribution Analysis
As discussed in Section 4.2, the effectiveness of training-free length extrapolation methods is influenced not only by the algorithm itself but also by the model’s understanding of positional intervals within its training context window. We propose a new comparison approach to fairly compare different length extrapolation methods while eliminating the positional interval understanding bias inherent to the model. Instead of applying extrapolation methods across a full context window, we first restrict them to a smaller positional interval range. We then extend this range to match the model’s training context window and compare the resulting attention distribution against the model’s original attention distribution. In this scenario, the closer the extrapolated attention distribution is to the model’s original attention distribution, the more effectively the extrapolation method performs within a smaller positional interval range, indicating that it aligns more closely with the model’s inherent understanding. Consequently, as the model’s understanding improves, the performance of the length extrapolation method will also improve. More concretely, we applied these training-free length extrapolation methods to Llama3-8b-ins-2k (Llama3-2k) and Llama3-4k backbones, and compared the resulting attention score distribution with Llama3-8k. The results, shown in Figure 4(a), indicate that GALI consistently demonstrates a significantly smaller gap compared to other methods, whether extrapolating from 2k positional intervals or 4k positional intervals to 8k. Remarkably, when using 2k positional intervals, GALI outperforms Dyn-NTK, NTK, and YaRN even when they use 4k positional intervals. Beyond global attention comparison, we also examined the row-wise differences in attention score entropy differences between Llama3-2k and Llama3-8k, as attention entropy has been identified as a key factor affecting model performance(Zhang et al., 2024; Farquhar et al., 2024). As shown in Figures 4(b) and 4(c), GALI consistently exhibits the smallest row-wise entropy difference from Llama3-8k among all methods, indicating fewer attention outliers compared to other methods. In summary, GALI preserves attention score distributions at both global and local levels, achieving optimal extrapolation performance. As backbone models continue to improve their understanding of positional intervals, GALI is expected to further enhance performance in downstream tasks.

### v2

@@ 3.2 Attention Logit Interpolation
To eliminate attention logit outliers and ensure robust extrapolation, GALI performs interpolation directly at the attention logit level, bypassing the need to compute position embeddings for unseen positional intervals. Unlike the methods that manipulate positional embeddings, which often produce unstable or extreme logit when extrapolated, GALI approximates logit via local linear interpolation and stabilizes them with Gaussian noise. This design draws on observations about the behavior of RoPE: while it encodes relative positions with oscillatory trigonometric functions, these functions become numerically unstable when extrapolated beyond pretrained ranges. Instead of applying RoPE embeddings to interpolated positions, GALI interpolates between known attention logit corresponding to valid pretrained positional intervals.
Concretely, for two tokens $\bm{x}_{m},\bm{x}_{n}\in\mathbb{R}^{l}$ at positions $m$ and $n$ (which may be floats due to interpolation), we define their positional interval as $r=m-n$. When $r$ is an integer, the corresponding attention logit is already trained and can be used directly. When $r$ is fractional, GALI linearly interpolates between the logit at $\lfloor r\rfloor$ and $\lceil r\rceil$, and introduces noise proportional to the positional interval to preserve oscillatory behavior:
$a(\bm{x}_{m},\bm{x}_{n},r)=a(\bm{x}_{m},\bm{x}_{n},\lfloor{r}\rfloor)-[a(\bm{x}_{m},\bm{x}_{n},\lfloor{r}\rfloor)$ (3)
$-a(\bm{x}_{m},\bm{x}_{n},\lceil{r}\rceil)]*({r}{\%}1)+\mathcal{N}(0,{\frac{r}{L_{tr}}}^{2})$
To enable efficient matrix operations with the positional interval matrix $R$ in the computation process (Figure 1), we employ an approximate implementation by substituting $r$ with $r=\lceil{m}\rceil-n$, as elaborated in the pseudo-code in Appendix D.
By avoiding embedding-level extrapolation and operating directly on logit, GALI eliminates outliers and achieves robust length extrapolation over long sequences in a training-free manner.
@@ 4 Experiments
We evaluate GALI on Llama3-8B-ins models across two task categories: real-world long-context tasks and long-context language modeling tasks. For comparison, we implement all published training-free length extrapolation methods, including NTKbloc97 (2023b), Dyn-NTKbloc97 (2023a), YARNPeng et al. (2023), SelfExtendJin et al. (2024), and ChunkLlamaAn et al. (2024b). Appendix B details data statistics.
4.1 Experiments Setup
Real-world long-context task: We evaluate GALI on two widely used long-context benchmarks, LongBenchBai et al. (2024) and L-EvalAn et al. (2024a). For LongBench, we use 16 English datasets, while for L-Eval, we focus on closed-ended groups. For consistency, we follow the official task prompt templates and truncation strategies from the respective benchmarks.
Long-context language modeling task: To evaluate GALI’s long-context language modeling capabilities, we use the test split of PG19Rae et al. (2019), an open-vocabulary language modeling benchmark derived from Project Gutenberg.
@@ 4.1 Experiments Setup
Table 1: Performance comparison on LongBench. The best result in each experiment is bolded. Results marked with * are reported by LongBench Jin et al. (2024). The number following each method denotes the target context window size (e.g., 16k represents 16 × 1024 tokens). "Original" refers to evaluations conducted using the backbone model in the left column. Additional results using the Llama2-7B-Chat-4K backbone are provided in Appendix C.1.
Backbone models and baseline methods: We use Llama3-8b-ins-4k (Llama3-4k) and Llama3-8b-ins-8k (Llama3-8k) as backbone models, where the number following each model indicates its initial context window size. We obtain Llama3-4k backbone via modifying its max_position_embedding parameter. We use shorter-versions of Llama3-8b-ins over other LLMs with shorter training context windows like LLama2 since it cannot fully understand all pretrained positional intervals, which limits GALI’s effectiveness in practice. The effective understanding range of LLMs is shorter than their training context window, as evidenced in Jin et al. (2024); Hsieh et al. (2024). For the baseline methods, we compare with NTKbloc97 (2023b), Dyn-NTKbloc97 (2023a), YaRNPeng et al. (2023) using huggingface implementation and SelfExtendJin et al. (2024), ChunkLlamaAn et al. (2024b) with their official implementation. They are all of the training-free length extrapolation methods up to now. The implementation details of these methods can be found in Appendix E.
@@ 4.2 Real-World long-context Task Results
The LongBench results (Table 1) highlight GALI’s strong average performance on the Llama3-8b-ins backbone series, surpassing both the 4k and 8k backbones as well as all other methods. Notably, (1) when using the Llama3-8k backbone with a 32k context window, GALI’s average score improves only slightly over the 16k setting (by 0.21), whereas other methods—except ChunkLlama—achieve much larger gains (often exceeding 1 point ); (2) using a 16k context window on the Llama3-4k backbone yields better results than the same context window on Llama3-8k. Again, GALI exhibits only minor gains in this setting, while other methods show substantially larger improvements. These two observations demonstrate that: (1) GALI achieves stable and superior performance without requiring input-length-specific tuning; and (2) under current LLM architectures, performing extrapolation within a narrower positional interval range leads to better results, even on short-context tasks.
First, we note that Figure 6(b) shows most LongBench samples are shorter than 16k tokens when tokenized by Llama3. However, NTK, Dyn-NTK, YaRN, and SelfExtend all benefit significantly from expanding the context window to 32k using the Llama3-8k backbone, especially on HotpotQA and Musique [^2] SelfExtend is highly sensitive to its hyperparameters, as noted in its GitHub. So, our reproduced results on Llama3-8k at 32k far exceed those reported in the original paper., whose most samples fall below 16k tokens. This highlights a key weakness of global scaling methods: they require tuning of the global scaling factor according to input length. Moreover, simply matching the target context window to the input length is not sufficient, because misalignment in positional mapping can still lead to significant performance drops. In contrast, GALI improves by 2.75 on the longest dataset NarrativeQA, when moving from 16k to 32k, while performance on sequences shorter than 16k remains nearly unchanged. This confirms that GALI can achieve stable and superior results without any input-length-specific tuning. Meanwhile, ChunkLlama exhibits only minor fluctuations. These results reflect the fundamental differences between approaches:
GALI maps inputs to the full range of positional intervals learned in pretraining, while NTK, Dyn-NTK, YaRN, and SelfExtend map inputs into a fixed range based on the input length and the global scaling factor (e.g., in SelfExtend, a larger group size leads to a narrower range). ChunkLlama determines its effective positional mapping through hyperparameters like chunk size and local window.
Consider HotpotQA and Musique: a 32k context window on the Llama3-8k maps 16k-length inputs to [0, 4096), while a 16k window maps to [0, 8192). For SelfExtend, this mapping is not fixed and depends on its hyperparameters, but typically compresses the range. Thus, even though a 16k window is sufficient to cover the input, NTK, Dyn-NTK, YaRN, and SelfExtend still benefit from using 32k, as it places more tokens in narrower positional ranges. This implies that these methods are sensitive to the scaling factor, making it difficult to determine optimal settings in practice. GALI, by contrast, achieves stable and superior performance without such tuning. Although ChunkLlama also avoids scaling, it suffers from degraded performance due to the loss of local positional information. Second, it is important to recognize that LLMs interpret positional intervals differently depending on their training context window, as observed in Hsieh et al. (2024). This explains why a 16k window on Llama3-4k outperforms the same 16k setting on Llama3-8k: since LLMs are trained via next-token prediction, they are more familiar with shorter positional intervals. As a result, all methods except ChunkLlama perform better with a 16k context window on Llama3-4k than with the same context window on Llama3-8k. The L-Eval results further support our analysis. As shown in Table 2, GALI achieves the highest average performance across most configurations, except when using a 32k context window on Llama3-8k. This is consistent with the LongBench pattern: while 32k on Llama3-8k performs slightly better than 16k, it still falls short of 16k on Llama3-4k, again reflecting how narrower positional ranges improve extrapolation performance. Since Llama3 better understands shorter intervals, methods that remap text into a smaller positional range gain an advantage at 32k, while GALI’s full-span reuse may be less aligned in this case. However, when we tested GALI with a 32k context window on Llama3-4k to force it to operate within the [0, 4096) interval, it once again achieved the best results. Figure 2(b) summarizes these performance trends. Additionally, GSM, QuALITY, and TOEFL, whose inputs remain below 8k with the Llama3 tokenizer, show consistent gains for SE, YaRN, and NTK over the base model, as shown in Appendix C.1. These results confirm that mapping into narrower, well-trained positional intervals benefits extrapolation, even for short context tasks. Experiments across LongBench and L-Eval support two key conclusions: (1) GALI avoids global scaling, requires no input-length-specific tuning, and achieves superior and stable performance through logit-level interpolation and greedy reuse of pretrained intervals; and(2) Training-free extrapolation methods benefit from using narrower positional intervals, compensating for LLMs’ uneven positional understanding.
Table 2: Performance comparison on L-Eval. The best results are bolded. The GSM, QuALITY, and TOEFL were excluded here since their sequence lengths remain below 8192 tokens when using the Llama3 tokenizer, making them unsuitable for long-context evaluation. Results for these three datasets, along with those for all datasets using the Llama2-7B-Chat-4K (Llama2-4k) backbone, are provided in Appendix C.1.
@@
Table 3: Performance on the PG19 dataset across varying target context window sizes.
@@ 4.3 Long Language Modeling Task Results
The language modeling results are shown in Table 3. Due to OOM, we cannot get ChunkLlama’s PPL results when setting the maximum position embedding to 32768. Except for NTK and DYN-NTK, all methods maintained a stable PPL without exploding. While low PPL does not guarantee better real-world task performance, an exploding PPL is a clear indicator of performance degradation in downstream tasks. Notably, GALI achieved the second-lowest PPL, demonstrating superior stability in length extrapolation. We tested PPL using a 16k contest window with Llama2-4k backbone. Please refer to the Appendix C.2.
@@
Figure 3: Differences in attention score metrics and row-wise entropy across methods compared to the original LLM. (a) represents the attention score matrix differences. The values show the sum of the absolute differences between the attention scores of the length interpolation methods and the original model. (b) and (c) show row-wise entropy differences. 2K indicates methods using [0, 2048) positional intervals, while 4k corresponds to [0, 4096). The figures are generated using a sample from NarrativeQA with a prefill length of 8091 tokens. More details on the attention score distributions can be found in Appendix C.3.
@@ 4.4 Attention Distribution Analysis
By analyzing the attention logit distribution of GALI, we observe its local linear interpolation eliminating attention logit outliers produces an interpolated attention distribution that closely matches the original, allowing it to preserve the model’s native behavior when extrapolating to longer inputs. This enables the extrapolation process to fully benefit from the model’s pretrained capabilities. We design a new experiment to demonstrate the advantage of GALI’s ability to avoid attention logit outliers and maintain an attention distribution that best aligns with the original model. That is, evaluate extrapolation methods while controlling for the model’s inherent positional interval bias. Instead of applying extrapolation across the full training context window, we first restrict each method to a narrower positional interval range. We then extend this range to match the model’s training context window and compare the resulting attention distributions to those produced by the original model. A closer match indicates better alignment with the model’s native positional understanding. Consequently, as the positional understanding of the model improves, the effectiveness of the extrapolation method also improves if its attention distribution remains faithful to the original.
More concretely, we apply various training-free extrapolation methods to Llama3-8b-ins-2k (Llama3-2k) and Llama3-4k, and compare their attention score distributions against that of Llama3-8k. As shown in Figure 3(a), GALI consistently yields the smallest distribution gap, whether extrapolating from 2k or 4k to 8k. Remarkably, GALI using 2k intervals outperforms Dyn-NTK, NTK, and YaRN using 4k intervals. In addition to the comparison of the global attention score, we further examine the differences in row-wise attention entropy between Llama3-2k and Llama3-8k, as attention entropy has been shown to strongly correlate with model performance Zhang et al. (2024); Farquhar et al. (2024). As illustrated in Figures 3(b) and 3(c), GALI achieves the smallest row-wise entropy differences among all methods, indicating fewer attention outliers and greater local stability.
In summary, GALI’s attention logit interpolation, eliminating attention logit outliers, preserves the attention score distribution at both global and local levels, which underpins its strong performance in length extrapolation. As backbone models continue to improve their understanding of positional intervals, the ability of GALI to maintain distributional alignment is expected to further enhance its downstream performance.